A projectile has an initial launch speed of 18.0 m/s. If it is desired that it strikes a target 31.0 m away at the same elevation, what should be the projection angle?


1
Expert's answer
2022-03-14T09:45:45-0400

The range

"R=\frac{v_0^2\sin2\theta}{g}"

Hence

"\sin2\theta=gR/v_0^2\\
=9.81*31.0/18.0^2=0.939"

"\theta=34.9^\circ"